# version 0.6

### v3

@@ 5.1 The Linear-Quadratic Regulator
The Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
@@
    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive semi-definite matrices. Thus, the cost functions are convex.
@@
The LQR is a special case of a Markov decision process which we have described in chapter 3, since the transitions are stochastic. However, naive application of value iteration, policy iteration, or linear programming do not apply, since there are infinitely many states and actions.
@@ 5.2 Optimal Solution of the LQR
Nevertheless, it can be shown that the optimal policy for an LQR is a linear policy, i.e. a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact together with a remarkable property of the value function is formally given in the following theorem. We state the theorem for a finite number of iterations, and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
Since ${R}\succcurlyeq 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succcurlyeq{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
@@
where in the last equality we used the fact that $\mathop{\mbox{\bf E}}[\mathbf{w}]=0$, and that the perturbation is independent of the state and control. We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$.
@@
If the matrices above are not full rank, we use the notation $M^{-1}$ for the Moore-Penrose pseudo inverse. Plugging this in the expression for the value function, we get
@@
where ${S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})}$ is positive semi-definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
@@ 5.3.1 Solving the DARE
In the 1-D case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case.
@@ 5.4 Robust Control
One of the most important and prominent generalizations is that of robust control, also known as $H_{\infty}$-control. The goal of H-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, we aim to solve:
@@
Using techniques from dynamic games, the following can be proven, see e.g [ZDG96b, BB08].
@@
For the $H_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
@@ 5.5 Bibliographic Remarks
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to miss-specification of system parameters. This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary.
@@ 5.6 Exercises
Part II Basics of Nonstochastic Control
Chapter 6 Policy classes for dynamical systems
In chapter 5, we studied the LQR problem, and saw that for linear dynamical systems, the optimal control takes a very special form: it is a linear function of the state. The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state. It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm which we analyzed in a previous chapter.
@@
    The transition mapping changes is not fixed and changes online.
@@
This motivates the current chapter which studies more general policies. We consider several options for control policies, as well as the relationships between them, in the context of continuous control. A control policy can take action according to several signals that may be available and observable as listed below.
  1. 1.
    State Action Control policies.
    The system state, denoted $\mathbf{x}_{t}$, if it is observable. Policies that map the state to action are called State-Feedback or State-Action policies, and are the most commonly considered in the literature. As we have seen in chapter 5, the optimal policy for LDS with stochastic noise is a linear state-feedback policy.
    More generally, a State-Action policy may depend on several previous states in a linear or non-linear manner. The most commonly considered policy class is a linear combination of previous states, but any (non-linear) mapping of the form
    $\mathbf{u}_{t}=\pi(\mathbf{x}_{t-1:t-h})=\pi(\mathbf{x}_{t-1},...,\mathbf{x}_{t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{x}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}},$
    can be considered. The parameter $h>0$ is the number of most recent states that the policy considers when computing the next action.
  2. 2.
    Disturbance Action Control policies.
    If the perturbation to the system, denoted $\mathbf{w}_{t}$, is known, Disturbance Action Control policies (DAC) choose an action based upon the current and potentially previous disturbances,
    $\mathbf{u}_{t}=\pi(\mathbf{w}_{t-1:t-h})\ ,\ \pi\in{\mathbb{R}}^{d_{\mathbf{w}}\times h}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
    For time-invariant linear dynamical systems, the two policies classes, linear DAC and linear SAC, are equivalent, since the states and disturbances are linearly related under DAC and SAC policies. We give a formal theorem that quantifies this relationship later in this chapter.
    However, for certain scenarios, the policy class of DAC is significantly better behaved than state-action control policies from a computational viewpoint. We shall explore efficient algorithms that learn optimal policies from this class in the next chapter.
  3. 3.
    Observation Action Control policies.
    In the case of dynamical systems that are partially observed and/or unknown, both policy classes above are infeasible, since the state and disturbance are similarly unknown. A natural policy parameterization is linear in the past observations. A special case of such policies is the class of Linear Dynamical Controllers, as stated below.
  4. 4.
    Linear Dynamical Control policies.
    A particularly useful policy class for partially observed linear dynamical systems is Linear Dynamical Controllers. These policies simulate an “internal linear dynamical system” to estimate the hidden state, and choose the action at any time step as a linear function of this estimate.
    A linear dynamic controller $\pi$ has parameters $\left(A_{\pi},{B}_{\pi},C_{\pi},{D}_{\pi}\right)$, and chooses a control at time $t$ according to
    $\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}+D_{\pi}\mathbf{x}_{t}\ ,\ \mathbf{s}_{t+1}=A_{\pi}\mathbf{s}_{t}+B_{\pi}\mathbf{x}_{t}.$
  5. 5.
    Disturbance Response Control policies.
    Another important class of policies for control with partial observability is known as the disturbance response controllers, or DRC. The motivation for this control policy class is computational: in later chapters we show efficient algorithms that learn policies within this class.
    The main signal used by the DRC policy class is called Nature’s y’s, denoted by $\mathbf{y}^{\mathbf{nat}}_{t}$. This signal is the would-be observation at time $t$ assuming all inputs to the system were zero from the beginning of time. Mathematically, DRC policies choose an action as a function of previous Nature’s y’s signals, i.e.,
    $\mathbf{u}_{t}=\pi(\mathbf{y}^{\mathbf{nat}}_{t:t-h})\ ,\ \pi:{\mathbb{R}}^{d_{\mathbf{y}}\times(h+1)}\mapsto{\mathbb{R}}^{d_{\mathbf{u}}}.$
6.1 Relating the power of policy classes
We motivated the consideration of nonstandard policy classes by the need to deviate from the classical reinforcement learning model of Markov Decision Processes. Another important consideration is computational; some policy classes we consider will yield efficient algorithms, whereas in standard representation none are known.
To be able to give rigorous statements on the performance of methods that optimize these new policy classes, it is crucial to be able to relate their representation power. Furthermore, as we consider adversarial perturbations, we shall compare policies in terms of how they perform in hindsight, rather than using any absolute metric of optimality. This motivates our main comparison metric as follows.
@@ 6.1 Relating the power of policy classes
We say that a class of policies $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following clause is satisfied: for every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that, simultaneously for all sequences of $T$ disturbances and cost functions, it holds that
@@
The proof of this lemma is left as an exercise to the reader (Exercise 6.4). We now proceed to study the approximation relationships between the policy classes we surveyed earlier.
6.2 Quantitative Comparison of Policy Classes for Time Varying Linear Dynamical Systems
In this section we formally relate the power of different complexity classes for the special case of time-varying linear dynamical systems. Recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics, and thus are of broad interest.
@@ 6.2 Quantitative Comparison of Policy Classes for Time Varying Linear Dynamical Systems
$\mathbf{x}_{t+1}$ $=A_{t}\mathbf{x}_{t}+B_{t}\mathbf{u}_{t}+\mathbf{w}_{t}.$
A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$. Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems, and in Figure 8.1 for partially observed systems.
Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear State-Action Control policies of increasing diameter for linear time invariant systems.
@@ 6.2.1 Linear Policies
The most elementary and well-studied policy classes for control of linear dynamical systems are linear policies, i.e. those that choose a control signal as a (fixed) linear function of the state. The most important subcase of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state. This was formalized in Definition 2.4. Formally,
@@
A linear controller $K\in{\mathbb{R}}^{d_{u}\times d_{x}}$ is a linear operator mapping state to control. We denote by $\Pi_{\gamma}^{LIN}$ the set of all linear controllers that are $\gamma$-stabilizing and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.
6.2.2 Linear Dynamic Controllers
The next policy class is particularly useful for linear dynamical systems with partial observation. Linear Dynamical Controllers simulate an “internal linear dynamical system” so to recover a hidden state, and play a linear function over this hidden state. The formal definition is given below.
Definition 6.4 (Linear Dynamic Controllers).
$\mathbf{u}_{t}^{\pi}=C_{\pi}\mathbf{s}_{t}\ ,\ \mathbf{s}_{t}=A_{\pi}\mathbf{s}_{t-1}+B_{\pi}\mathbf{x}_{t}.$
Clearly the class of LDCs approximates linear policies without any error, since we can take $A_{\pi},B_{\pi},C_{\pi}$ to be zero, and obtain a purely linear policy.
We denote by $\Pi^{LDC}_{\gamma}$ the class of all LDC that are $\gamma$-stabilizing as per Definition 2.4 and Exercise 4.6. This implies the existence of parameters $\frac{\kappa}{\delta}\leq\gamma$ that satisfy
LDCs are an important policy class for control of linear dynamical systems since they perform favorably in presence of partial observability. Furthermore, the optimal control law for latent LDS with Gaussian disturbances and quadratic loss functions is contained in the class of LDCs.
6.2.3 Generalized Linear Controllers
We present a generalization of linear controllers below. This generalization class is at least as expressive as LDCs.
Definition 6.5 (Generalized Linear Controllers).
A generalized linear controller $\pi$ has parameters $M_{0:h}$, and chooses a control at time $t$ according to
$\mathbf{u}_{t}^{\pi}=\sum_{i=0}^{h}M_{i}\mathbf{x}_{t-i}.$
An important concept for generalized linear controllers is that of lifting, or creating a larger linear system that subsumes the GLC. Given a certain GLC with memory $h$, consider the system given by
$\mathbf{z}_{t+1}=\tilde{A}_{t}\mathbf{z}_{t}+\tilde{B}_{t}\mathbf{u}_{t}+\tilde{C}_{t}\mathbf{w}_{t},$
where
$\tilde{A}_{t}=\begin{array}[]{|c|c|c|c|c|}\hline\cr A_{t}&0&\ldots&0&0\\ \hline\cr I&0&\ldots&0&0\\ \hline\cr 0&I&\ldots&0&0\\ \hline\cr\vdots&\vdots&\cdots&\vdots&\vdots\\ \hline\cr 0&0&\cdots&I&0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{B}_{t}=\begin{array}[]{|c|}\hline\cr B_{t}\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}\ \ ,\ \ \tilde{C_{t}}=\begin{array}[]{|c|}\hline\cr I\\ \hline\cr 0\\ \hline\cr 0\\ \hline\cr\vdots\\ \hline\cr 0\\ \hline\cr\end{array}.$
It can be seen that a GLC for the original system is simply a linear controller $\mathbf{u}_{t}^{\pi}=[M_{0}\dots M_{h}]\mathbf{z}_{t}$ for the lifted system! This observation will be useful to us later on.
We denote by $\Pi^{GLC}_{h,\gamma}$ the class of all GLC with $\sum_{i=0}^{h}\|M_{i}\|\leq\gamma$, and $\mathbf{M}_{1:h}$ that is $\gamma$-stabilizing as per Definition 2.4. The following lemma (proof left as as exercise) shows that GLC approximate LDC (and of course linear controllers):
Lemma 6.6.
The class $\Pi^{GLC}_{h,\gamma}$ $\varepsilon$-approximates the class $\Pi^{LDC}_{\gamma^{\prime}}$ for $h=\Omega(\gamma^{\prime}\log(\frac{\gamma^{\prime}}{\varepsilon}))$, and $\gamma=2\gamma^{\prime 3}$.
The class of GLC is thus very general and intuitive. It is, however, difficult to work with for the regret metric, due to its counterfactual structure.

### v4

@@ 5.1 The Linear Quadratic Regulator
The (time-invariant) Linear Quadratic Regulator (LQR) input is a linear dynamical system that evolves according to
@@
    $Q\in{\mathbb{R}}^{d_{x}\times d_{x}}$, ${R}\in{\mathbb{R}}^{d_{u}\times d_{u}}$ are positive definite matrices. Thus, the cost functions are convex.
@@
The LQR is a special case of a Markov decision process which we have described in chapter 3, since the transitions are stochastic. However, we cannot naive apply value iteration, policy iteration, or linear programming since there are infinitely many states and actions.
@@ 5.2 Optimal Solution of the LQR
Nevertheless, it can be shown that the optimal policy for an LQR is a linear policy, i.e. a linear mapping from ${\mathbb{R}}^{d_{x}}$ to ${\mathbb{R}}^{d_{u}}$! This astonishing fact, together with a remarkable property of the value function, is formally given in the following theorem. We state the theorem for a finite number of iterations, and for a linear time-invariant system. Generalizing this to a time-varying linear system is left as exercise 5.6.
@@
Since ${R}\succ 0$, a minimum occurs at $\mathbf{u}=0$. Hence, ${S}_{T}={Q}\succ{0}$, ${K}_{T}={0}$ and $c_{T}=0$.
@@
where in the last equality we used the fact that $\mathop{\mbox{\bf E}}[\mathbf{w}]=0$, and that the perturbation is independent of the state and control. We henceforth denote the second moment of the noise by $\mathop{\mbox{\bf E}}[\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}I$. This implies that for any matrix $M$ which is independent of $\mathbf{w}$, we have $\mathop{\mbox{\bf E}}[\mathbf{w}^{\top}M\mathbf{w}]=\mathop{\mbox{\bf E}}[M\mathbf{w}\mathbf{w}^{\top}]=\sigma^{2}\cdot M\bullet I=\sigma^{2}{\bf Tr}(M)$.
@@
Notice that the matrix $R+B^{\top}S_{t}B$ is full rank since $R\succ 0$. Plugging this in the expression for the value function, we get
@@
where ${S_{t-1}=Q+K_{t-1}^{\intercal}RK_{t-1}+(A-BK_{t-1})^{\intercal}S_{t}(A-BK_{t-1})}$ is positive definite, and $c_{t-1}=c_{t}+\sigma^{2}{\bf Tr}(S_{t})$.
@@ 5.3.1 Solving the DARE
In the 1-D case, although the right side of the equation is a rational function, we can multiply both sides by ${R+B^{\top}SB}$ to form a quadratic equation in ${S}$. Solving the DARE therefore corresponds to finding a non-negative root of a quadratic polynomial in the one-dimensional case. The reason we only consider the non-negative root is since the value function is non-negative by our choice of quadratic convex costs.
@@ 5.4 Robust Control
One of the most important and prominent generalizations is that of robust control, also known as $\text{H}_{\infty}$-control. The goal of $\text{H}_{\infty}$-control is to come up with a more robust solution, by considering the optimal solution vs. any possible perturbation sequence from a certain family. Formally, we aim to solve:
@@
Using techniques from dynamic games, the following can be shown. References are provided in the bibliographic section.
@@
For the $\text{H}_{\infty}$ control problem with quadratic convex loss functions, the optimal solution is a linear policy of the state given by
@@ 5.5 Bibliographic Remarks
The development of $H_{\infty}$ control was driven by an observation made in the seminal paper of Doyle [Doy78] that, unlike in the fully observed case, the solution to the optimal control problem for partially observed linear systems (which we study in later chapters) is not naturally robust to miss-specification of system parameters. This suggests that robustness specifications constitute yet an another axis of performance for control algorithms, and may at times be at odds with the objective of optimal control. For in-depth treatment of $H_{\infty}$ control, see [ZDG96b]. The text [BB08] presents a succinct derivation of robust controllers entirely in the state space, as the solution to a dynamic game between the controller and an adversary, see also [Tu99] for a recent exposition.
@@ 5.6 Exercises
5.3. Extend the proof of Theorem 5.2 to positive semi-definite $Q,R$, rather than positive definite matrices.
5.4. Prove Theorem 5.4 as a dynamic two-player zero-sum game, where a unique saddle point solution exists. Hint: you can formulate this problem as a min-max optimization problem within a finite horizon.
5.5. Prove that the function $f(\mathbf{x})=\mathbf{x}^{\top}M\mathbf{x}$ is convex if and only if $M\succeq 0$ is positive semi-definite.
Part II Basics of Online Control
Chapter 6 Regret in Control
In the first part of this book and chapter 5 we studied linear dynamical systems with quadratic costs, namely the LQR problem. We saw that in this case the optimal control takes a very special form: it is a linear function of the state. The Markovian property of the Markov Decision Process framework for reinforcement learning suggests that the optimal policy is always a function of the current state. It is characterized by the Bellman optimality equation, and can be computed in complexity proportional to the number of states and actions using, for example, the value iteration algorithm which we analyzed in a previous chapter.
@@
    The transition mapping is not fixed and changes online.
@@
This motivates the current chapter which considers regret analysis as a computational tractable alternative to optimal or robust control.
We start by considering regret in the decision making framework of online convex optimization. We then consider several classes of control policies, as well as the relationships between them, in the context of regret analysis.
6.1 Online Convex Optimization
In this section we give a brief introduction to the theory of Online Convex Optimization. The reader is referred to the bibliographic section for a comprehensive survey.
$\mathbf{x}_{t}^{\mathcal{A}}={\mathcal{A}}(f_{1},...,f_{t-1})\in\mathcal{K}.$
We define the regret of ${\mathcal{A}}$ after $T$ iterations as:
$\mathrm{{regret}}_{T}({\mathcal{A}})=\sup_{\{f_{1},...,f_{T}\}\subseteq{\mathcal{F}}}\left\{\sum_{t=1}^{T}f_{t}(\mathbf{x}_{t}^{\mathcal{A}})-\min_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})\right\}.$ (6.1)
6.1.1 Online Gradient Descent
Perhaps the simplest algorithm that applies to the most general setting of online convex optimization is Online Gradient Descent (OGD). Pseudo-code is given in Algorithm 3.
Algorithm 3 Online Gradient Descent
2:  for $t=1$ to $T$ do
3:     Play $\mathbf{x}_{t}$ and observe cost $f_{t}(\mathbf{x}_{t})$.
4:     Update and project:
$\mathbf{y}_{t+1}=\mathbf{x}_{t}-\eta_{t}\nabla f_{t}(\mathbf{x}_{t})$
$\mathbf{x}_{t+1}=\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}_{t+1}\|^{2}=\prod_{\mathcal{K}}[\mathbf{y}_{t+1}]$
5:  end for
  • $D$ = Euclidean diameter of the decision set, i.e.
    $\max_{\mathbf{x},\mathbf{y}\in\mathcal{K}}\|\mathbf{x}-\mathbf{y}\|.$
    $\forall t\ ,\ \|\nabla f_{t}(\mathbf{x}_{t})\|\leq G.$
Theorem 6.1.
Proof.
Let $\mathbf{x}^{\star}\in\operatorname*{\arg\min}_{\mathbf{x}\in\mathcal{K}}\sum_{t=1}^{T}f_{t}(\mathbf{x})$. Define $\nabla_{t}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}\nabla f_{t}(\mathbf{x}_{t})$. By convexity
$f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})\leq\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$ (6.2)
$\|\mathbf{y}_{t+1}-\mathbf{x}^{\star}\|^{2}\ =\ \left\|\mathop{\Pi}_{\mathcal{K}}(\mathbf{x}_{t}-\eta_{t}\nabla_{t})-\mathbf{x}^{\star}\right\|^{2}\leq\left\|\mathbf{x}_{t}-\eta_{t}\nabla_{t}-\mathbf{x}^{\star}\right\|^{2}.$ (6.3)
Hence,
$2\nabla_{t}^{\top}(\mathbf{x}_{t}-\mathbf{x}^{\star})$$\leq$ $\ \frac{\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}-\|\mathbf{x}_{t+1}-\mathbf{x}^{\star}\|^{2}}{\eta_{t}}+\eta_{t}G^{2}.$ (6.4)
Summing (6.2) and (6.1.1) from $t=1$ to $T$, and setting $\eta_{t}=\frac{D}{G\sqrt{t}}$ (with $\frac{1}{\eta_{0}}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}0$):
$\leq\sum_{t=1}^{T}\|\mathbf{x}_{t}-\mathbf{x}^{\star}\|^{2}\left(\frac{1}{\eta_{t}}-\frac{1}{\eta_{t-1}}\right)+G^{2}\sum_{t=1}^{T}\eta_{t}$ $\frac{1}{\eta_{0}}\stackrel{{\scriptstyle\text{\tiny def}}}{{=}}0,$
$\|\mathbf{x}_{T+1}-\mathbf{x}^{*}\|^{2}\geq 0$
$\leq D^{2}\frac{1}{\eta_{T}}+G^{2}\sum_{t=1}^{T}\eta_{t}$ telescoping series
$\leq 3DG\sqrt{T}.$
6.2 Regret for Control
The most important consideration to explore beyond classical models of control, such as optimal and robust control, is computational. Even in linear control with convex loss functions, the optimal policy can be complex to describe, see exercises section.
A prominent approach to deal with computational difficulty is to consider a less stringent criteria for optimality. Instead of the optimal policy, we can consider all policies in a certain policy class, and try to compete with the optimal policy within the class in terms of regret.
As alluded to earlier, the most natural classes of policies are computationally intractable. We thus resort to learning new classes that are computationally favorable. But to relate the performance of our methods, we require a language of comparing different policy classes. It is crucial to be able to relate the representation power of different policy classes. This motivates our main comparison metric as follows.
@@ 6.2 Regret for Control
We say that a class of policies $\Pi_{1}$ $\varepsilon$-approximates class $\Pi_{2}$ if the following is satisfied: for every $\pi_{2}\in\Pi_{2}$, there exists $\pi_{1}\in\Pi_{1}$ such that for all sequences of $T$ disturbances and cost functions, it holds that
@@
Proof.
Recall that for policy class $\Pi$ and algorithm $\mathcal{A}$,
$\mathrm{{regret}}_{T}(\mathcal{A},\Pi)=\max_{\mathbf{w}_{1:T}:\|\mathbf{w}_{t}\|\leq 1}\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi\in\Pi}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi},\mathbf{u}_{t}^{\pi})\right)$
It follows from definition 6.2 that
$\min_{\pi_{2}\in\Pi_{2}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})$ $=\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}^{*}},\mathbf{u}_{t}^{\pi_{2}^{*}})$ $\pi_{2}^{*}=\operatorname*{\arg\min}_{\pi\in\Pi_{2}}\sum_{t}c_{t}(\mathbf{x}^{\pi}_{t},\mathbf{u}_{t}^{\pi})$
$\geq\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T$ Definition 6.2
$\geq\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T$
Thus, for any fixed disturbance sequence $\mathbf{w}_{1:T}$,
$\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi_{2}\in\Pi_{2}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{2}},\mathbf{u}_{t}^{\pi_{2}})$
$\leq\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\left(\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})-\varepsilon T\right)$
$=\left(\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})-\min_{\pi_{1}\in\Pi_{1}}\sum_{t=1}^{T}c_{t}(\mathbf{x}_{t}^{\pi_{1}},\mathbf{u}_{t}^{\pi_{1}})\right)+\varepsilon T$
Since this holds for any fixed disturbance sequence $\mathbf{w}_{1:T}$, it also holds for the maximum over all disturbance sequences $\mathbf{w}_{1:T}$ with $||\mathbf{w}_{t}||\leq 1$. ∎
We now proceed to define and study approximation relationships between different policy classes.
6.3 Expressivity of Control Policy Classes
In this section we formally relate the power of different complexity classes for linear time-invariant linear dynamical systems. The definitions and results can be extended to time-varying linear dynamical systems, and recall from Chapter 4 that time varying linear dynamical systems were been proposed by Lyapunov as a general methodology to study nonlinear dynamics.
@@ 6.3 Expressivity of Control Policy Classes
$\mathbf{x}_{t+1}$ $=A\mathbf{x}_{t}+B\mathbf{u}_{t}+\mathbf{w}_{t}.$
A sequence of stabilizing linear controllers for such a system is denoted by $\{K_{t}\}$. Schematically, the relationships we describe below are given in Figure 6.1 for fully observed systems. In later chapters of this text we study policies for partially observed systems.
Figure 6.1: Schematic relationship between linear Disturbance-Action Control policies and linear state-action policies of increasing diameter for linear time invariant systems.
@@ 6.3.1 Linear Policies
The most basic and well-studied policy classes for control of linear dynamical systems are linear policies, i.e. those that choose a control signal as a (fixed) linear function of the state. The most important subcase of linear policies are those that are $\gamma$-stabilizing, meaning that the signal they create does not cause an unbounded state. This was formalized in Definition 2.4. Formally,
@@
A linear controller $K\in{\mathbb{R}}^{d_{x}\times d_{u}}$ is a linear operator mapping state to control. We denote by $\Pi_{\gamma}^{L}$ the set of all linear controllers that are $\gamma$-stabilizing according to definition 2.4 and are bounded in Fronbenius norm by $\|K\|\leq\gamma$.